Please help me with this question and help me understand that.

Be sure to answer all parts. The dissociation of molecular iodine into iodine atoms is represented as I2(g) - 21g) At 1000 K, the equilibrium constant Kc for the reaction is 3.80 × 10-5. Suppose you start with 0.0458 mol of I2 in a 2.33-L flask at 1000 K. What are the concentrations of the gases at equilibrium? What is the equilibrium concentration of I2? What is the equilibrium concentration of I?
Answer #1

The given chemical reaction is

I_2 _{(g)} \rightleftharpoons 2I_{(g)}

Equilibrium constant K_c of a reaction is given by

K_c = \frac{\prod [Product]^a}{\prod [Reactant]^b}

Where a and b are the stoichiometric coefficient in the balanced reaction.

Hence, K_c for our reaction is written as

K_c = \frac{[I]^2}{I_2} = 3.80 \times 10^{-5}

where the concentrations are their equilibrium values.

The initial amount of I2 is 0.0458 mol.

Volume of the flask = 2.33 L

Hence, initial concentraion of the I2 gas is

\frac{0.0458 \ mol}{2.33 \ L} \approx 0.020 \ M

Each mole of I_2 dissociates to give 2 moles of I .

Hence, we can write the ICE(Initial, Change, Equilibrium) chart as follows:

[I_2] [I]
Initial 0.020 0
Change -x +2x
Equilibrium 0.020-x 2x

Hence, using the dissociation constant equation

K_c = \frac{[I]^2}{I_2} = 3.80 \times 10^{-5} \\ \Rightarrow 3.80 \times 10^{-5} = \frac{(2x)^2}{0.020-x} \\ \Rightarrow x = - 4.4 \times 10^{-4} \ or \ x = 4.31 \times 10^{-4}

Since, x is a concentration and cannot be negative, x = 4.31 \times 10^{-4} \ M .

Equilibrium concentration of I_2 = 0.020 M -x.

But since x << [I_2] , we can say that the concentration of I_2 does not change.

Hence,

[I_2]_{(eq)} = 2.0 \times 10^{-2} \ M - 4.31 \times 10^{-4} \ M \approx 2.0 \times 10^{-2} \ M

Now,

[I]_{eq} = 2x = 2 \times 4.31 \times 10^{-4} \ M = 8.62 \times 10^{-4} \ M

Hence, equilibrium concentration of I_2 is 0.020 M.

Equilibrium concentraion of I is 0.000862 M.